# MEV Capture Through Time-Advantaged Arbitrage

### v1

@@ Abstract
As blockchains begin processing significant economic activity, the ability to include and order transactions inevitably becomes highly valuable, a concept known as Maximal Extractable Value (MEV). This makes effective mechanisms for transaction inclusion and ordering, and thereby the extraction of MEV, a key aspect of blockchain design. Beyond traditional approaches such as ordering in a first-come-first-serve manner or using priority fees, a recent proposal suggests auctioning off a time advantage for transaction inclusion. In this paper, we investigate this time advantage mechanism, focusing specifically on arbitrage opportunities on Automated Market Makers (AMMs), one of the largest sources of MEV today. We analyze the optimal strategy for a time-advantaged arbitrageur and compare the profits generated by various MEV extraction methods. Finally, we explore how AMMs can be adapted in the time advantage setting to capture a portion of the MEV.
@@
Optimization MEV Arbitrage AMMs Latency.
@@ 1 Introduction
Decentralized Finance (DeFi)has played a central role in the Ethereum ecosystem ever since its inception, encompassing numerous active users and many of the most successful applications. More recently, DeFihas been expanding to Layer-2 (L2)scaling solutions. Trading volumes on L2sare rising consistently, with some rollups already executing a higher number of daily swaps than the Ethereum network [4, 8].
Two main features of rollups drive this trend in DeFi. First, rollups offload most of the computation required to verify and execute transactions outside the Ethereum network, only storing condensed batches of transactions on Ethereum. This allows them to run more efficiently and reduce transaction costs for end users. Second, rollups store the condensed batches of transactions and proofs on Ethereum, thus inheriting Ethereum’s security guarantees.
As DeFiactivity on rollups grows, so do the incentives for various opportunistic economic practices collectively known as Maximal-Extractable Value (MEV). MEVis a phenomenon first described for the Ethereum blockchain [3], and corresponds to the value that can be extracted from the inclusion and reordering of transactions in a block. Currently, the most popular forms of MEVon Ethereum are arbitrages (that try to capitalize on price discrepancies between different exchanges), liquidations (where collateral from debt repayment in lending protocols can be purchased at a discount), and sandwiching (which involves wrapping a victim’s swap transaction between two new transactions in such a way that provides a worse trade execution to the victim while tracking a small profit) [12]. On some of the most prominent rollups today, arbitrages and liquidity occur in similar volumes to Ethereum, while sandwich attacks have not been observed [14].
@@ 1.1 MEV Capture Mechanisms
With both DeFiand arbitrage activity increasing, policies to include and order transactions have become a key design choice for rollups. Initially, most rollups employed a simple design to order transactions solely based on when the centralized sequencer received them. This ordering system is known as First-Come First-Serve (FCFS), and both optimistic (such Arbitrum) and ZK (such as ZKsync) rollups still use it. In this system, transaction fees are determined by the state transition function, and users cannot pay a premium for faster inclusion. Instead, users who value fast transaction inclusion are motivated to invest in latency infrastructure in order to improve their latency to the sequencer.
@@
As expected, the transaction ordering policy significantly influences how the profit from MEVextraction is distributed among participants. For instance, on Ethereum, most profits are paid to validators through MEV-Boost, while on Arbitrum and for FCFS, more generally, profits stay with MEVextractors and are likely invested in improving their access to sequencers (e.g., through latency improvements). Avoiding such latency competitions would arguably lead to a more efficient market design. This can be achieved by letting extractors bid for the opportunity to extract MEVin an auction. A recent proposal by Arbitrum [1] involves introducing an auction market to sell a time advantage, with the proceeds going to the rollup protocol. Concretely, users bid for the right to access an “express lane” where transactions are sequenced immediately. This access is sold to a single user for a fixed time interval. Transactions submitted by the remaining users are artificially delayed by the sequencer for a predefined time, resulting in the user who purchases the time advantage being able to guarantee their transactions being sequenced ahead of its competitors.
@@ 1.2 Contributions
  • We simulate the expected profits an arbitrageur could extract in some key liquidity pools under three separate sequencing regimes — FCFS, PGA, and Timeboost. We find that the theoretical finding in the previous step, that it is always optimal to wait with the trade as long as possible, is sometimes invalidated. A potential explanation of this finding is that future price distribution is not independent of the past price movement, which can actually be seen as an outside market imperfection.
@@ 2 Model Description
We study the scenario in which one single actor has a time advantage of $T_{w}$ over everybody else when submitting transactions to a rollup (or blockchain) sequencer over a time period of $T$. Practically, the time advantage is achieved by delaying all transactions by other actors by $T_{w}$. To assess the value of this advantage, we examine the potential arbitrage profits that can be extracted with this time edge.
@@
Specifically, we examine the profit potential from performing arbitrage trades between an AMMpool on the rollup and an external venue on which the same token pair is traded. The external venue could be thought of as comprising the entire outside market. We assume the price impact of trading on the external venue is negligible compared to the price impact on the AMM.  [^1] This is the case when the external venue is assumed to have infinitely deep liquidity, or at least significantly deeper liquidity than the AMMpool.
Consider an AMMpool with a trading fee of $f$ that facilitates trades between two assets — a risky asset and a numéraire. We assume that the risky asset’s external market price changes over time, while the price on the AMMis only changed by arbitrage trades. In other words, we assume that the price discovery for the risky asset takes place on the external market. [^2] Furthermore, we ignore uninformed (non-arbitrage, “noise”) trades on the AMM. Let $p$ denote the price of the risky asset on the external market relative to the price on the AMM. For instance, $p=0.9$ indicates that the external price is 10% lower than the AMMprice. Initially, the prices on the AMMand the external market prices are assumed to be the same, that is, $p=1.0$.
The actor with the time advantage can monitor price discrepancies between the external market and the AMM. When an arbitrage opportunity arises, they can decide whether and when to capitalize on it. Realizing the arbitrage profit involves both a trade on the external market and a trade on the AMM. For the time-advantage arbitrageur, the AMMtrade will immediately reach the AMM. Although other arbitrageurs can also detect these opportunities, their AMMtrades will be delayed. As a result, there exists a time window of length $T_{w}$ during which no other arbitrageur can interact with the AMMas a reaction to the arbitrage opportunity, since their transactions are subject to this delay.
@@
Additionally, we assume that arbitrage opportunities that are not exploited within a time of $T_{w}$ by the time-advantaged arbitrageur will be exploited by other arbitrageurs. Note that this is a conservative assumption with respect to the profits of the time-advantaged arbitrageur. [^3] It is not guaranteed that delayed arbitrageurs will always attempt to exploit arbitrage opportunities due to the risk of partly execution and changes of the external price during the delay period. Importantly, any strategy employed by the time-advantaged arbitrageur under these assumptions will remain effective regardless of the actions of the delayed arbitrageurs.
@@ 2.1 The Time-Advantaged Arbitrageur
The time-advantaged arbitrageur decides whether to execute an arbitrage trade or wait based on the current price difference $p$, the elapsed time $t_{w}$ within their advantage window, and the elapsed time $t$ within the overall period $T$. We aim to determine the optimal strategy for this decision and the resulting profits.
We model the overall time horizon $T$ as a series of discrete time steps of size $\Delta$. At each time step — i.e., at time $t=0,\Delta,2\Delta,\ldots$ – the arbitrageurs, particularly the one with a time advantage, have the opportunity to submit trades to both the AMMand the external market.
Given the relative price difference $p$ between the AMMand the external market, we consider the maximal arbitrage trade, i.e., the arbitrage trade that yields the maximal profit, and denote the profits from executing it as $profit(p)$. This profit is measured in the numéraire and is considered relative to the AMMpool’s value, which is also measured in the numéraire at the initial AMMprice. Note that if the price difference is too small, no profitable arbitrage trade will be possible due to the trading fees on the AMMand, in this case, $profit(p)=0$.
Due to the AMM’s trading fee, the maximal arbitrage does not completely eliminate the price difference. Let $pArb(p)$ represent the remaining price difference after the maximal arbitrage trade is executed at the price difference $p$. Finally, let $poolVal(p)$ denote the relative change in the pool’s value resulting from the maximal arbitrage trade executed at the price difference $p$.
Let $opt(p,t_{w},t)$ represent the expected profit a time-advantaged arbitrageur can achieve between time $t$ and $T$ by employing the optimal strategy, given a price difference of $p$ at time $t$ and an elapsed time $t_{w}$ within the advantage window. That is,
@@
Note that this definition of $opt()$ assumes that the external price of the risky asset follows a memoryless process, which means that future price changes are independent of past prices. Consequently, $opt()$ depends solely on the current price difference $p$ rather than the entire history of past prices.
@@
We also assume that taking any potential arbitrage opportunity at the end of the delay time $T_{w}$ is optimal. When doing so, the arbitrageur instantly makes $profit(p)$ and resets their advantage window $t_{w}=0$. Furthermore, the price difference is moved to $pArb(p)$, which means that the expected later profits between time $t$ and $T$ are given by $opt(pArb(p),0,t)$. However, the later profits are measured relative to the pool value after the arbitrage. So, this term needs to be multiplied by the relative change in the pool values. Altogether, this means that:
@@
For any other point in time, that is, when $t_{w}\neq T_{w}$ and $t\neq T$, the arbitrageur will choose between making an arbitrage trade at $t$ or waiting until the next time step $t+\Delta$. Their optimal decision will be the one that, in expectation, promises more future profit.
In the case of waiting, the expected future profit corresponds to the expected future profit of the next time step, assuming no change in the AMMpool price (since no arbitrage trade is being made). If we define $p_{\Delta}$ as the random variable that describes the relative price change over the time step $\Delta$, then the expected profit of waiting is given by
@@
Note that the expectation is taken over the random variable $p_{\Delta}$.
@@
Therefore, the expected profit of the optimal strategy for the advantaged arbitrageur is defined as the maximum of the expected profit of these two possible decisions:
@@
Finally, we want to express $opt(pArb(p),0,t)$ in terms of results from later times steps. This is needed for the empirical analysis later on. Since no profitable arbitrage trade is possible by definition of $pArb(p)$, the only option is to wait for a time step $\Delta$. Therefore,
@@ 3.1 Arbitrage Profits for Constant Product Pools
If we assume that the pool being arbitraged is a constant product market maker (CPMM), we can represent the profit function, ($profit(p)$) and the remaining price difference and pool size after the arbitrage trade ($pArb(p)$ and $poolVal(p)$ respectively) in terms of the price difference $p$ and the trading fee $f$:
$pArb(p)=\begin{cases}\frac{1}{1-f}&\text{if }p>\frac{1}{1-f}\\ p&\text{if }1-f<p<\frac{1}{1-f}\\ 1-f&\text{if }p<1-f\end{cases}$ (1)
$profit(p)=\begin{cases}\frac{1}{2}p-\frac{1}{\sqrt{1-f}}\sqrt{p}+\frac{1}{2}\frac{1}{1-f}&\text{if }p>\frac{1}{1-f}\\ 0&\text{if }1-f<p<\frac{1}{1-f}\\ \frac{1}{2}\frac{1}{1-f}p-\frac{1}{\sqrt{1-f}}\sqrt{p}+\frac{1}{2}&\text{if }p<1-f\end{cases}$ (2)
$poolVal(p)=\sqrt{\frac{p}{pArb(p)}}=\begin{cases}\sqrt{1-f}\sqrt{p}&\text{if }p>\frac{1}{1-f}\\ 1&\text{if }1-f<p<\frac{1}{1-f}\\ \frac{1}{\sqrt{1-f}}\sqrt{p}&\text{if }p<1-f.\end{cases}$ (3)
For conciseness, we provide the full derivation of these formulas in Section 0.A of the appendix.
@@ 3.2 Analytical Results
In particular, the latter holds when prices follow a geometric Brownian motion — the standard model for asset prices. There, the price changes $p_{\Delta}$ are distributed independently and log-normally, i.e.,

### v2

@@ Abstract
As blockchains begin to process significant economic activity, the ability to include and order transactions inevitably becomes highly valuable, a concept known as Maximal Extractable Value (MEV). This makes effective mechanisms for transaction inclusion and ordering, and thereby the extraction of MEV, a key aspect of blockchain design. Beyond traditional approaches such as ordering in a first-come-first-serve manner or using priority fees, a recent proposal suggests auctioning off a time advantage for transaction inclusion. In this paper, we investigate this time advantage mechanism, focusing specifically on arbitrage opportunities on Automated Market Makers (AMMs), one of the largest sources of MEV today. We analyze the optimal strategy for a time-advantaged arbitrageur and compare the profits generated by various MEV extraction methods. Finally, we explore how AMMs can be adapted in the time advantage setting to capture a portion of the MEV.
@@
MEV Arbitrage AMMs Latency Rollups.
@@ 1 Introduction
Decentralized Finance (DeFi)has played a central role in the Ethereum ecosystem ever since its inception, encompassing many of its most-used applications. DeFi encompasses a range of financial services, including token trading, as well as borrowing and lending, which are facilitated through smart contracts deployed on the blockchain. More recently, DeFihas begun expanding to Layer-2 (L2)scaling solutions. Trading volumes on L2sare rising consistently, with some rollups already processing more token swaps than the Ethereum itself [5, 9].
Two main features of rollups drive this trend in DeFi. First, rollups offload most of the computation required to verify and execute transactions outside the Ethereum network, only storing condensed batches of transactions on Ethereum. This allows them to run more efficiently and reduce transaction costs for end users. Second, rollups store the condensed batches of transactions and proofs on Ethereum, thus inheriting (some of) Ethereum’s security guarantees.
As DeFiactivity on rollups grows, so do the incentives for various opportunistic economic practices collectively known as Maximal-Extractable Value (MEV). MEV– a phenomenon first described for the Ethereum blockchain [4] – describes the value that can be extracted from the ability to include and reorder transactions on a blockchain. Currently, the most popular forms of MEVare arbitrages (that try to capitalize on price discrepancies between different on-chain trading venues), liquidations (where collateral from debt repayment in lending protocols can be purchased at a discount), and sandwiching (which involves wrapping a victim’s swap transaction between two new transactions in such a way that provides a worse trade execution to the victim while making a small profit) [13]. On the most prominent rollups today, arbitrages and liquidity occur in similar volumes to Ethereum, while sandwich attacks have not been observed [15].
@@ 1.1 MEV Capture Mechanisms
With both DeFiand arbitrage activity increasing, policies to include and order transactions have become a key design choice for rollups. Initially, most rollups employed a simple design to order transactions solely based on when the centralized sequencer received them. This ordering system is known as First-Come First-Serve (FCFS), and both optimistic (such as Arbitrum) and ZK (such as ZKsync) rollups still use it. In this system, transaction fees are determined by the state transition function, and users cannot pay a premium for faster inclusion. Instead, users who value fast transaction inclusion are motivated to invest in latency infrastructure in order to improve their latency to the sequencer.
@@
As expected, the transaction ordering policy significantly influences how the profit from MEVextraction is distributed among participants. For instance, on Ethereum, most profits are paid to validators through MEV-Boost, while on Arbitrum and for FCFS, more generally, profits stay with MEVextractors and are likely invested in improving their access to sequencers (e.g., through latency improvements). Avoiding such latency competitions would arguably lead to a more efficient market design. This can be achieved by letting extractors bid for the opportunity to extract MEVin an auction. A recent proposal by Arbitrum [2] involves introducing an auction market to sell a time advantage, with the profits going to the rollup protocol. Concretely, users bid for the right to access an “express lane” where transactions are sequenced immediately. This access is sold to a single user for a fixed time interval. Transactions submitted by the remaining users are artificially delayed by the sequencer for a predefined time, resulting in the user who purchases the time advantage being able to guarantee their transactions being sequenced ahead of its competitors.
@@ 1.2 Contributions
  • We simulate the expected profits an arbitrageur could extract in some key liquidity pools under three separate sequencing regimes – FCFS, PGA, and Timeboost. We find that the theoretical finding in the previous step, that it is always optimal to wait with the trade as long as possible, is sometimes invalidated. A potential explanation of this finding is that future price distribution is not independent of the past price movement, which can actually be seen as an outside market imperfection.
@@ 2 Model Description
We study the scenario in which one single actor has a time advantage of $T_{w}$ over everybody else when submitting transactions to a rollup (or blockchain) sequencer over a time period of length $T$. This time advantage is achieved by delaying the processing of all other actors’ transactions by $T_{W}$, while the advantaged actor’s transactions are sequenced immediately. In other words, one actor purchases access to a “fast lane”, allowing their transactions to bypass an artificially applied delay of $T_{w}$ during the regular processing of transactions. Apart from the applied delay, transactions are ordered on a first-come, first-served basis. Note that there is no limit on the number of transactions that can be submitted through the fast lane during the period $T$ of access.
@@
To assess the value of this advantage, we examine the potential arbitrage profits that can be extracted with this time edge. Specifically, we examine the profit potential from performing arbitrage trades between an AMMpool on the rollup and an external venue on which the same token pair is traded. The external venue could be thought of as comprising the entire outside market. We assume the price impact of trading on the external venue is negligible compared to the price impact on the AMM.  [^1] This is the case when the external venue is assumed to have infinitely deep liquidity, or at least significantly deeper liquidity than the AMMpool.
2.1 Arbitrage Opportunities on AMMs
We consider an AMMpool with a trading fee $f$, facilitating trades between two assets. These pools collect token deposits from LPsas reserves, against which traders can swap for a fee. Exchange rates are algorithmically determined – e.g., by maintaining a constant product of the reserves, in the simplest case [1].
We consider a pool containing a risky asset and a numéraire, and assume the risky asset’s price fluctuates on the outside market over time. The AMM price adjusts solely through arbitrage trades with the outside market, reflecting our assumption that price discovery occurs externally. We ignore uninformed ("noise") trades on the AMM.
Let $p$ represent the price of the risky asset on the external market relative to its price on the AMM. Specifically, for an AMM price $P_{0}$ and an external market price $P$, we define $p=P/P_{0}$. For example, $p=0.9$ indicates that the external price is 10% lower than the AMMprice. Initially, we assume the prices on the AMMand the external market are equal, i.e., $p=1.0$.
Arbitrageurs monitor price discrepancies between the external market and the AMM. When an arbitrage opportunity arises, they compete to exploit it. Realizing arbitrage profits involves executing one trade each on the external market and the AMM. Notably, only the first arbitrageur to interact with the AMMcan perform a profitable arbitrage, as their trade moves the AMM pool price towards the external market price.
In the setting we analyze, the time-advantaged arbitrageur’s trades are immediately processed by the AMM, while transactions from other arbitrageurs are delayed. Consequently, a time window of length $T_{w}$ exists, starting when an arbitrage opportunity emerges, during which no other arbitrageur can respond to the opportunity due to the imposed delay. This allows the time-advantaged arbitrageur to decide whether and when to capitalize on the opportunity during this time window, contrasting with a pure first-come, first-serve setting where competition forces immediate extraction.
@@ 2.1 Arbitrage Opportunities on AMMs
Additionally, we assume that any arbitrage opportunities not exploited by the time-advantaged arbitrageur within a time of $T_{w}$ of their emergence will be seized by other arbitrageurs. Note that this is a conservative assumption with respect to the profits of the time-advantaged arbitrageur. [^2] It is not guaranteed that delayed arbitrageurs will always attempt to exploit arbitrage opportunities due to the risk of partly execution and changes of the external price during the delay period. Importantly, any strategy employed by the time-advantaged arbitrageur under these assumptions will remain effective regardless of the actions of the delayed arbitrageurs.
@@ 2.2 The Time-Advantaged Arbitrageur
[…]
We model the overall time horizon $T$ as a series of discrete time steps of size $\Delta$. At each time step – i.e., at time $t=0,\Delta,2\Delta,\ldots$ – the arbitrageurs, particularly the one with a time advantage, have the opportunity to submit trades to both the AMMand the external market.
Given the relative price difference $p$ between the AMMand the external market, we consider the maximal arbitrage trade, i.e., the arbitrage trade that yields the maximal profit, and denote the profits from executing it as $profit(p)$. This profit is measured in the numéraire and is considered relative to the AMMpool’s value, which is also measured in the numéraire at the AMMprice before the AMM trade. Note that if the price difference is too small, no profitable arbitrage trade will be possible due to the trading fees on the AMMand, in this case, $profit(p)=0$. The profit function for a constant product AMM pool is derived in Section 3.1 and plotted in Figures 1.
Due to the AMM’s trading fee, the maximal arbitrage does not completely eliminate the price difference. Let $pArb(p)$ represent the remaining relative price difference after the maximal arbitrage trade is executed at the relative price difference $p$. Finally, let $poolVal(p)$ denote the relative change in the pool’s value resulting from the maximal arbitrage trade executed at the relative price difference $p$.
Let $opt(p,t_{w},t)$ represent the expected profit a time-advantaged arbitrageur can achieve between time $t$ and $T$ by employing the optimal strategy, given a relative price difference of $p$ at time $t$ and an elapsed time $t_{w}$ within the advantage window. That is,
@@
Note that this definition of $opt()$ assumes that the external price of the risky asset follows a memoryless process, which means that future price changes are independent of past prices. Consequently, $opt()$ depends solely on the current relative price difference $p$ rather than the entire history of past prices.
@@
Furthermore, we assume that it is always optimal for the arbitrageur to exploit any available arbitrage opportunity at the end of the delay time $T_{w}$ (since the opportunity will otherwise be taken by other arbitrageurs). When doing so, the arbitrageur instantly makes a profit of $profit(p)$ and resets their advantage window $t_{w}=0$ (since the other arbitrageurs need to react to the change of the AMM price). Additionally, the arbitrageur expects further future profits between time $t$ and $T$. Since, the arbitrage trade has moved the relative price difference to $pArb(p)$, the expected future profits are given by $opt(pArb(p),0,t)$. However, the future profits are measured relative to the pool value after the arbitrage. So, this term needs to be multiplied by the relative change in the pool value $poolVal(p)$. Combining immediate profits and scaled future profits, the total expected profit is:
@@
For any other point in time, that is, when $t_{w}\neq T_{w}$ and $t\neq T$, the arbitrageur must decide between two actions:
  1. 1.
    Execute an arbitrage trade immediately at time $t$.
  2. 2.
    Wait until the next time step $t+\Delta$.
The arbitrageur will chose the action which yields the higher expected future profits.
If the arbitrageur chooses to wait, the expected future profit depends on the expected relative price difference at the next time step. Let $p_{\Delta}$ be the random variable describing the relative price change over the time interval $\Delta$. The expected profit of waiting is then calculated as the profit from the optimal strategy at the next step, given the updated price $p\cdot p_{\Delta}$ and the updated times $t_{w}+\Delta$ and $t+\Delta$:
@@
Here, the expectation is taken over the random variable $p_{\Delta}$.
@@
Therefore, the advantaged arbitrageur’s optimal expected profit is given by the maximum of the expected profit of their two possible decisions:
@@
Finally, to obtain a recursive definition, we express $opt(pArb(p),0,t)$ in terms of results from times steps later than $t$. This is needed for the empirical analysis later on. Since, by definition of $pArb(p)$, no profitable arbitrage trade is possible in this situation, the arbitrageur’s only option is to wait for a time step $\Delta$. Therefore,
@@ 3.1 Arbitrage Profits for Constant Product Pools
If the pool being arbitraged is a constant product market maker (CPMM), the profit function ($profit(p)$), and the remaining relative price difference and pool size after the arbitrage trade ($pArb(p)$ and $poolVal(p)$, respectively) in terms of the relative price difference $p$ and the trading fee $f$ is:
$profit(p)=\begin{cases}\frac{1}{2}p-\frac{1}{\sqrt{1-f}}\sqrt{p}+\frac{1}{2}\frac{1}{1-f}&\text{if }p>\frac{1}{1-f}\\ 0&\text{if }1-f\leq p\leq\frac{1}{1-f}\\ \frac{1}{2}\frac{1}{1-f}p-\frac{1}{\sqrt{1-f}}\sqrt{p}+\frac{1}{2}&\text{if }p<1-f\end{cases}$ (1)
$pArb(p)=\begin{cases}\frac{1}{1-f}&\text{if }p>\frac{1}{1-f}\\ p&\text{if }1-f\leq p\leq\frac{1}{1-f}\\ 1-f&\text{if }p<1-f\end{cases}$ (2)
$poolVal(p)=\sqrt{\frac{p}{pArb(p)}}=\begin{cases}\sqrt{1-f}\sqrt{p}&\text{if }p>\frac{1}{1-f}\\ 1&\text{if }1-f\leq p\leq\frac{1}{1-f}\\ \frac{1}{\sqrt{1-f}}\sqrt{p}&\text{if }p<1-f.\end{cases}$ (3)
For conciseness, we provide the full derivation of these formulas in Section 0.A of the appendix. Figure 1 plots the profit function for a trading fee of $f=0.05\%$.
@@ 3.2 Analytical Results
In particular, the latter holds when prices follow a geometric Brownian motion – the standard model for asset prices. There, the price changes $p_{\Delta}$ are distributed independently and log-normally, i.e.,